OpenStudy (anonymous):

wo functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 2 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x).

OpenStudy (anonymous):

wo functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 2 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x).

OpenStudy (sohailiftikhar):

it's x=5 you can find this by putting =5 in both functions

OpenStudy (sohailiftikhar):

so answer is 5 at which f(x)=g(x)

OpenStudy (anonymous):

LOL YO THANKS U HELP ME SO MUCH ON THIS TEST

OpenStudy (sohailiftikhar):

lol np

iYuko (iyuko):

@geny55

OpenStudy (sohailiftikhar):

what ?????